OpenStudy (ggrree):

I'm feeling a bit rusty with probabiliy. Could somebody quickly explain to me how to use permutations and combinations? I understand what they mean, but I'm confused at what the xPy or the xCy notation means( where x and y are numbers.

OpenStudy (anonymous):

\[ \large xPy = ^xP_y = \frac{x!}{(x-y)!} \] \[ \large xCy = ^xC_y = \frac{x!}{y! \times (x-y)!} \] where \( x,y \in \mathbb{N} \)

OpenStudy (anonymous):

xPy: It means that there are (x) objects going into (y) categories, order matters (ie) (x) different cars competing for (y) places in a race: xPy xCy: It means there are (x) objects going into (y) categories, but order DOES NOT matter: ie) (x) different cars competing for (y) identical parking spots

OpenStudy (anonymous):

Just remember "Permutation" is counting the different ways something can "Permeate" so order matters. "Combination" is counting different ways to combine the same things so order does not matter.

OpenStudy (anonymous):

right ...switched this two

OpenStudy (anonymous):